# Querying Autonomous Vehicle Point Clouds: Enhanced by 3D Object Counting with CounterNet

### v1

@@ E.3. Model Selection
(1) Compare results of using 6 models and 9 models. Table 18 and Figure 15 show the results of using 6 models for model selection. Compared to the results with 9 models, we removed the dominant models from the 9-model set (i.e., $CN_{pt=4}^{o=0}$, $CN_{pt=4}^{o=0.1}$, and $CN_{pt=4}^{o=0.2}$). According to Table 18, after removing the $CN_{pt=4}$ series models, the new dominant model without adjustment becomes $CN_{pt=9}$. However, with the adjustment, the dominant model shifts to $CN_{pt=9}^{o=0.1}$, while $CN_{pt=2}^{o=0.1}$ remains the second dominant model. From the results in Figure 15, it is evident that applying adjustments dramatically improves accuracy. When comparing the adjusted results to the best-performing model ($CN_{pt=4}^{o=0.2}$), even the weaker-performing models achieve comparable results through model selection with adjustments.
@@ E.4. Ablation Studies
Dynamic Thresholding. As explained in Section 5.1, we proposed using dynamic thresholding to filter out noise for peak detection. We compared the results with and without dynamic thresholding for the model CN-p4o. In our evaluation, we selected three fixed thresholds: 0.5, 0.52, and 0.54, for comparison. As shown in Figure 17, using a fixed threshold resulted in significant deviations across object categories. For instance, the motorcycle and bicycle categories experienced extremely low performance with thresholds of 0.5 and 0.52, while the car and truck categories, with a threshold of 0.54 performed poorly. In contrast, dynamic thresholding provided a more balanced performance across all categories, achieving consistently good results.
Weighted loss. In Section 5.1, we introduced the weighted loss during the model training with partitions. We introduce the weighted loss intent to improve the prediction with the frame with more objects appear. By refer to the data statistics of nuScenes dataset in Appendix E.1, we can see that the weighted loss contributed more to the low frequency objects, such as truck, trailer and construction vehicles. While for the high frequency appearance objects, such as car, pedestrian, and barrier, the improvement with partition is limited.
@@ F.1. Categorical Query Results
As shown in Table 14 and Table 4, the performance improvement of our proposed CounterNet ($CN$) on the KITTI dataset is less pronounced compared to the improvements observed on the nuScenes dataset when compared to the detection model CenterPoint (CP). For the SELECT-COUNT operation, while the performance improvements for the ”car” and ”cyclist” categories are not substantial, a noticeable improvement is observed for the ”pedestrian” category.
Regarding partitioning and overlap, the results for our proposed solution are consistent with those observed on the nuScenes dataset. However, considering that the scene scale of nuScenes is approximately four times larger than that of KITTI, partitioning has a more negative impact on performance in the KITTI dataset. Nevertheless, the use of overlap partially mitigates this negative effect.
F.1.1. Model Selection
@@
Table 20 and Figure 19 present the results of model selection with six models on the KITTI dataset. Overall, the findings are consistent with those observed on the nuScenes dataset. As shown in Table 20, after applying the adjustment, the frames are primarily concentrated on four partitions with a 0.2 overlap rate ($CN_{pt=4}^{o=0.2}$) and two partitions with a 0.2 overlap rate ($CN_{pt=2}^{o=0.2}$). Furthermore, as illustrated in Figure 19, similar to the nuScenes results, the model selection yields balanced performance across all categories.

### v2

@@ E.3. Model Selection
(1) Comparison results of using 6 models and 9 models. Table 18 and Figure 15 show the results of using 6 models for model selection. Compared to the results with 9 models, we remove the dominant models from the 9-model set (i.e., $CN_{pt=4}^{o=0}$, $CN_{pt=4}^{o=0.1}$, and $CN_{pt=4}^{o=0.2}$). According to Table 18, after removing the $CN_{pt=4}$ series models, the new dominant model without adjustment becomes $CN_{pt=9}$. However, with the adjustment, the dominant model shifts to $CN_{pt=9}^{o=0.1}$, while $CN_{pt=2}^{o=0.1}$ remains the second dominant model. From the results in Figure 15, it is evident that applying adjustments dramatically improves accuracy. When comparing the adjusted results to the best-performing model ($CN_{pt=4}^{o=0.2}$), even the weaker-performing models achieve comparable results through model selection with adjustments.
@@ E.4. Ablation Studies
Dynamic Thresholding. As explained in Section 5.1, we propose using dynamic thresholding to filter out noise for peak detection. We compare the results with and without dynamic thresholding for the model CN-p4o. In our evaluation, we select three fixed thresholds: 0.5, 0.52, and 0.54, for comparison. As shown in Figure 17, using a fixed threshold results in significant deviations across object categories. For instance, the motorcycle and bicycle categories experience extremely low performance with thresholds of 0.5 and 0.52, while the car and truck categories, with a threshold of 0.54, perform poorly. In contrast, dynamic thresholding provide a more balanced performance across all categories, achieving consistently good results.
Weighted loss. In Section 5.1, we introduce the weighted loss used during model training with partitions. This weighted loss is intended to improve predictions on frames where more objects appear. By referring to the data statistics of the nuScenes dataset in Appendix E.1, we observe that the weighted loss contributes more significantly to low-frequency object categories, such as trucks, trailers, and construction vehicles. In contrast, for high-frequency object categories—such as cars, pedestrians, and barriers—the improvement from partitioning is limited.
@@ F.1. Categorical Query Results
As shown in Table 14 and Table 4, the performance improvement of our proposed CounterNet ($CN$) on the KITTI dataset is less pronounced compared to the improvements observed on the nuScenes dataset when evaluated against the detection model CenterPoint (CP). For the SELECT-COUNT operation, while the performance gains for the ”car” and ”cyclist” categories are modest, a noticeable improvement is observed for the ”pedestrian” category.
Regarding partitioning and overlap, the results for our proposed method are consistent with those observed on the nuScenes dataset. However, since the scene scale in nuScenes is approximately four times larger than that of KITTI, partitioning has a more negative impact on performance in the KITTI dataset. Nevertheless, the use of overlap partially mitigates this adverse effect.
Model Selection. Table 20 and Figure 19 present the results of model selection with six models on the KITTI dataset. Overall, the findings are consistent with those observed on the nuScenes dataset. As shown in Table 20, after applying the adjustment, the frames are primarily concentrated on four partitions with a 0.2 overlap rate ($CN_{pt=4}^{o=0.2}$) and two partitions with a 0.2 overlap rate ($CN_{pt=2}^{o=0.2}$). Furthermore, as illustrated in Figure 19, similar to the nuScenes results, the model selection yields balanced performance across all categories.